OpenStudy (anonymous):

the point (1, -square root of 3) is on the terminal side of an angle in standard position. Give the smallest positive angle measure in both degrees and radians.

OpenStudy (watchmath):

\( \tan \theta = \frac{-\sqrt{3}}{1}\) Whaat is \(\theta\) ?

OpenStudy (anonymous):

thats wut im tryin to find out

OpenStudy (watchmath):

Remember \(\sin 60^\circ = \sqrt{3}/2\) and \(\cos 60^\circ =1/2\).